OpenStudy (anonymous):

How do I solve 8x^9 over 3x^7

OpenStudy (anonymous):

x^9/x^7 = x^(9-7) = x^2 so you have (8x^2)/3

OpenStudy (anonymous):

Okay Thanks

OpenStudy (dan815):

u know how powers work

OpenStudy (anonymous):

separate the problem into parts part 1 = 8/3 part 2 = x^9 / x^7 First part stays the same but the second can be simplified. Rules of exponents says when dividing subtract. so x^9 / x^7 becomes x^2. Final answer= 8x^2 / 3

OpenStudy (anonymous):

@dan815 not really

OpenStudy (dan815):

okay well x^7 basically means x*x*x*x**x*x*x 7 times

OpenStudy (anonymous):

and when we multiply do we add??? @engstat

OpenStudy (anonymous):

well yeah I know that @dan815

OpenStudy (dan815):

okay so lets think about it

OpenStudy (dan815):

if you multiply x 9 times on the top and on on the bottom you are multiplying x 7 times

OpenStudy (dan815):